OpenStudy (anonymous):

A movie theater charges $6.50 for matinee showings and $8.75 for evening showings. Yesterday the theater sold 378 tickets for a total revenue of $2,929.50. How many matinee tickets were sold?

OpenStudy (anonymous):

x+y=378 So x=378-y and then 6.5x+8.75y=6.5(378-y)+8.75y=2457-6.5y+… 2929.5=2457+2.25y 2.25y=472.5 okay so your answer would be y=210

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

np

OpenStudy (anonymous):

ONE MORE QUESTION

OpenStudy (anonymous):

How many solutions does the system Test Image and Test Image have? Explain.

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

how Many Solutions does the system y=-7x+3 and y+7x=10 have? explain

terenzreignz (terenzreignz):

Is this a quiz?

OpenStudy (anonymous):

yes